# Global survey of star clusters in the Milky Way VIII. Cluster formation and evolution

### v1

@@ 6 Comparison to a cluster sample based on Gaia DR3
As mentioned in Sect. 2, mass determinations of large cluster catalogues based on Gaia astrometry have so far been relatively sparse in the literature. Due to the high scientific interest in the Gaia data releases (Gaia Collaboration et al. 2016a, 2018, 2021, 2023), the state of the literature on Milky Way clusters using Gaia data has been rapidly evolving, so this will likely change in the near future. It seems therefore expedient to demonstrate that the CAMF and CAMF models can prove useful tools for analysing cluster systems also in the context of Gaia-based cluster catalogues.
@@
As an in-depth analysis of the catalogue’s completeness does not yet exist, we use the simple mass-based completeness model that was presented in Hunt & Reffert (2024) as an approximate estimate.
@@
Using the given completeness distances, ages, masses and corresponding uncertainties from the catalogue, we can construct CAMFs as we did using MWSC data. However, the $40\mathrm{M}_{\odot}$ lower mass cutoff enforces different mass limits than for the MWSC.
We obtain a total surface density of
@@
We observe that like for the MWSC sample, there is evidence for a period of enhanced mass loss at early ages from the change in the high-mass contours, as well as some effective upper age limit above which cluster densities drop rapidly.
The major difference is that both enhanced mass loss and age limit, as well as the cluster density maximum, appear at earlier ages than in the MWSC sample, with a constant high mass edge for $\tau<7.3$ and a high-age edge in the range of $\tau=$ 9–9.5. The density maximum occurs at $\tau\approx 8$ in contrast to $\tau\approx 8.7$ for the MWSC sample.
The high-age edge is further more strongly dependent on mass as well as more diffuse. As we will discuss below, this may be related to the age-mass uncertainties and their dependence on cluster age.
In order to quantitatively compare the modelled with the observed CAMF, we again need to take the uncertainties in age and mass into account. We find that age uncertainties vary weakly for ages $\tau=$ 7–9, typically remaining within $0.23\pm 0.06$ dex. For $\tau<7$, they are slightly reduced. No clear dependence on mass is apparent. In the same age range, mass uncertainties show little systematic variation with age apart from a slight increase for $\tau<7$, but a linear decrease of about 0.025 dex per decade in mass, with typical values at $0.07\pm 0.03$. As the cluster sample spans less than three decades in mass, this trend can account for less than 40% of the total variance in mass uncertainties. For $\tau>9$, age uncertainties, and to a lesser degree mass uncertainties, are higher and vary more strongly, with age uncertainties at $0.30\pm 0.19$ dex and mass uncertainties at $0.08\pm 0.04$ dex.
The fact that this increase in uncertainty coincides with the age limit and corresponding drop in the number of clusters observed may suggest that some of these clusters are poorly parametrised ‘outliers’ in the sense that they are strongly observationally scattered from a different part of the true CAMF. The generally higher age uncertainties may partially explain the more diffuse high-age edge of the CAMF compared to the MWSC sample.
We here use the simplest possible model of constant age and mass errors with
@@
while acknowledging that this is not accurate at high ages. However, at the level of this proof-of-concept analysis, we deem it sufficient.
Using this error model in conjunction with our CAMF model, we can now compare our best-fitting model for MWSC data to the Gaia-based CAMF, rescaling the CFR parameter $\beta$ such that the total cluster density for $\mu>1.6$ matches the observation. This results in an alternate choice of $\beta=3.037\mathrm{kpc^{-2}Myr^{-1}}$, about three times higher than for the MWSC sample.
As we noted above, the observed CAMF generally sits at lower ages compared to the MWSC data. Likewise, the CAMF of the best-fitting model has high-age edge, density maximum and violent relaxation mass loss at higher ages compared to observation.
Using steps of 0.05 dex, we find that a shift of the model to lower ages by 0.8 dex minimises the KLD between observed CAMF and error-smoothed model CAMF to $D_{\mathrm{KL}}=0.113$ from $D_{\mathrm{KL}}=0.666$ without an age shift. Such a shift in cluster age is equivalent to rescaling the parameters $\beta$, $c$, and $t_{\mathrm{V}}$, which respectively govern cluster formation rate, lifetimes, and the violent relaxation time scale.
@@
We finally use the KLD and the above error model to perform a model fit using MCMC sampling as we did for the MWSC sample. For simplicity, we reuse the same parameter constraints as they fit also this sample. For the Gaussian priors, we adopt the same means as a baseline, but use twice the standard deviations to account for the Gaussian priors being specific to the MWSC sample and avoid unduly constraining the parameter space.
The MCMC takes 221000 steps with an average acceptance rate of 0.161 until convergence. With a burn-in of 4404 samples and a thinning factor of 528, the final posterior sample contains 21320 parameter sets. The best model has a KLD of $D_{\mathrm{KL}}=0.034$, which represents a significant improvement compared to the age-shifted best-fitting MWSC model. The rightmost column of Table 4 contains a list of this model’s parameters with uncertainties derived from the posterior distribution.
@@
Indeed the best model, with its CAMF shown in Fig. 12, does not feature a break in the cluster lifetime-mass relation and has a high value of $a_{1}$ at 0.538, resulting in slow dissolution at the end of clusters’ lifetimes and a high number of clusters with masses below $40\mathrm{M}_{\odot}$.
The Schechter cutoff mass at $\log m_{\mathrm{S}}/\mathrm{M}_{\odot}=4.743$ corresponds to lifetimes close to the age limit and seems mostly sufficient to reproduce it.
@@
The model contours in the left panel of Fig. 12 appear to match the observed CAMF fairly well.
Synthetic cluster samples drawn from the model contain on average $2447\pm 41$ clusters for a total surface density of $\Sigma_{\mu>1.6}=257.22\pm 0.50\mathrm{kpc}^{-2}$ above the mass limit, which appears consistent with observations within the stochastic variation. However, all 5000 synthetic samples have lower values of the KLD than the observed sample, indicating the presence of systematic differences between observation and model that cannot be explained by stochastic noise.
@@ 7 Summary and conclusions
In this paper we investigated the parameter space of a theoretical model for the cluster age-mass function (CAMF), which was introduced in Just et al. (2023), where a base model was presented. The model is constructed with a constant cluster formation rate (CFR), a two-power law cluster initial mass function (CIMF) with an exponential Schechter cutoff, and a cluster bound mass function $m(M,t)$ describing the cluster mass loss as function of initial mass and age. The cluster mass evolution is adapted to the models of Shukirgaliyev et al. (2017, 2018) based on a centrally concentrated star formation efficiency.
@@
A posterior analysis by random sampling of the models was used to determine whether a configuration similar to the observed cluster sample was a likely configuration of such a model sample. For this, the different statistics as well as the total cluster number were used to quantify deviations between models and their samples, and assess whether the observed cluster sample could be differentiated from a model sample based on the observed values of the different statistics. The base model, published in Just et al. (2023), could be differentiated from observations by $\chi^{2}$, KLD, and total cluster number, but not by the maximum likelihood method. The models returned from $\chi^{2}$ and maximum likelihood could be differentiated less easily, but also showed deviations in at least two of the statistics considered. The model returned from the KLD in particular was best able to reproduce the observed cluster number and could only be differentiated from observations by the maximum likelihood method.
Using principal component analysis, we derived the ellipsoidal shape of the 11-dimensional posterior distribution in parameter space for the different MCMC runs. The semi-major axes of the ellipsoids (given by the square root of the normalised covariance matrix eigenvalues) vary by a factor of four, e.g. from ${\sqrt{\lambda_{0}}=0.341}$ to ${\sqrt{\lambda_{10}}=1.486}$ for the $D_{\mathrm{KL}}$-fit (see also Fig. 5). These eigenvalues give the variance of the parameter vector in the direction of the eigenvector relative to the individual parameter variances. An eigenvalue smaller than 1 corresponds to a combination of parameters that is more tightly constrained than the parameters are individually, while an eigenvalue larger than 1 corresponds to a combination that is constrained less so. The eigenvectors corresponding to both the smallest and largest eigenvalues are similar for the posteriors of the different runs, indicating that the different fit statistics realise the same constraints imposed by the observed CAMF. The model parameters are very well constrained in two dimensions fixing the upper age limit and the position of the maximum of the CAMF and badly constrained in one dimension corresponding to low-mass end of the distribution.
@@
The base model was constructed to reproduce the mass functions in four age bins (Just et al. 2023). The focus was on the younger ages $\log t/\mathrm{yr}<8.3$ and the maximum of the number surface density. The best model resulting from the $D_{\mathrm{KL}}$-fit, which we adopt as our best-fitting model, provides a similarly good representation of these features of the CAMF and a significantly improved model for the old and the high-mass regions.
@@
However, there are two main challenges in inferring model parameters and selecting best-fit models we identified over the course of this work. Both challenges are expected to not be specific to the model functions and the cluster data we used, but intrinsic to the problem of modelling the evolution of a cluster system. We identified these first in the context of the MWSC dataset and found them to also be relevant when using the Hunt & Reffert (2023, 2024) catalogue. The first challenge lies in the strongly correlated nature of model parameter space, which makes constraining individual parameters difficult. For instance, the rates of cluster formation and dissolution must be balanced such that the model reproduces the observed cluster content, but individually each of these rates are badly constrained. A possible remedy of this problem may lie in better and more sophisticated constraints on cluster mass evolution from prior physics considerations such as N-body simulations. The second challenge lies in the measuring of observation-model correspondence, where uncertainties in the measurements of observed cluster properties, the completeness limits of the observed cluster sample and the varying cluster density across the age-mass plane must be taken into consideration. In particular, the resolution of the observed CAMF in the age-mass plane is limited by both sample size and measurement uncertainties, which can make approaches that work without explicit construction of the observed CAMF, such as our expected log-likelihood approach, more attractive. We demonstrated the use of forward modelling to check the behaviour of a fit statistic for a given model, and discussed how a model fitting procedure can be limited by the fit statistic used. More work is needed to identify efficient statistics for measuring the correspondence of a CAMF model to a cluster sample, especially in view of the extensive, high-quality cluster catalogues based on Gaia DR3 (Gaia Collaboration et al. 2023) that are currently emerging (e.g. Hunt & Reffert 2023).
@@ Appendix A Rapid mass evolution in the observed CAMF using linear age binning
We have argued in Sect. 2 that the constant high-mass edge for $\tau<7.8$ in the observed CAMF is evidence for rapid mass evolution during the violent relaxation phase.
One way to more accessibly visualise this is to remove the effect of logarithmic age binning by normalising using linear bin sizes in age instead (as also used in Fig. 8 to visualise the factorisation of the CAMF in initial-mass).
@@ Appendix B The modelled CAF
As discussed in Sect. 1, the CAF is typically of interest in the context of fitting cluster evolution models. While our 2D approach does not consider or fit the CAF explicitly, correspondence between observed and modelled CAF should nonetheless arise by construction.
For our model, the CAF $\eta(t)$, giving the cluster surface density per age interval, can be computed at some age $t$ by integrating the CAMF over all masses.
Recalling Eqs. (10) and (46), with $M_{0}(t)$ as inverse of $t_{\mathrm{Cl}}(M)$, we find

### v2

@@ 6 Comparison to a cluster sample based on Gaia DR3
As we mention in Sect. 2, mass determinations of large cluster catalogues based on Gaia astrometry have so far been relatively sparse in the literature. Due to the high scientific interest in the Gaia data releases (Gaia Collaboration et al. 2016a, 2018, 2021, 2023), the state of the literature on Milky Way clusters using Gaia data has been rapidly evolving, so this will likely change in the near future. It seems therefore expedient to demonstrate that the CAMF and CAMF models can prove useful tools for analysing cluster systems also in the context of Gaia-based cluster catalogues.
@@
As an in-depth analysis of the catalogue’s completeness does not yet exist, we used the simple mass-based completeness model that is presented in Hunt & Reffert (2024) as an approximate estimate.
@@
Using the given completeness distances, ages, masses and corresponding uncertainties from the catalogue, we constructed CAMFs as we did using MWSC data. However, the $40\mathrm{M}_{\odot}$ lower mass cutoff enforced different mass limits than for the MWSC.
We obtained a total surface density of
@@
We observe that like for the MWSC sample, there is evidence for a period of enhanced mass loss at early ages from the change in the high-mass contours, as well as some effective upper age limit above which cluster densities drop rapidly. The major difference is that both enhanced mass loss and age limit, as well as the cluster density maximum, appear at earlier ages than in the MWSC sample, with a constant high mass edge for $\tau<7.3$ and a high-age edge in the range of $\tau=$ 9–9.5. The density maximum occurs at $\tau\approx 8$ in contrast to $\tau\approx 8.7$ for the MWSC sample. The high-age edge is further more strongly dependent on mass as well as more diffuse. As we will discuss below, this may be related to the age-mass uncertainties and their dependence on cluster age.
In order to quantitatively compare the modelled with the observed CAMF, we again needed to take the uncertainties in age and mass into account. We find that age uncertainties vary weakly for ages $\tau=$ 7–9, typically remaining within $0.23\pm 0.06$ dex. For $\tau<7$, they are slightly reduced. No clear dependence on mass was apparent. In the same age range, mass uncertainties show little systematic variation with age apart from a slight increase for $\tau<7$, but a linear decrease of about 0.025 dex per decade in mass, with typical values at $0.07\pm 0.03$. As the cluster sample spans less than three decades in mass, this trend can account for less than 40% of the total variance in mass uncertainties. For $\tau>9$, age uncertainties, and to a lesser degree mass uncertainties, are higher and vary more strongly, with age uncertainties at $0.30\pm 0.19$ dex and mass uncertainties at $0.08\pm 0.04$ dex.
The increase in uncertainty coincides with the age limit and corresponding drop in the number of clusters observed. This may suggest that some of these clusters are poorly parametrised ‘outliers’ in the sense that they are strongly observationally scattered from a different part of the true CAMF. The generally higher age uncertainties may partially explain the more diffuse high-age edge of the CAMF compared to the MWSC sample.
We here used the simplest possible model of constant age and mass errors with
@@
while acknowledging that this is not accurate at high ages. However, at the level of this proof-of-concept analysis, we deemed it sufficient.
Using this error model in conjunction with our CAMF model, we compared our best-fitting model for MWSC data to the Gaia-based CAMF, rescaling the CFR parameter $\beta$ such that the total cluster density for $\mu>1.6$ matches the observation. This resulted in an alternate choice of $\beta=3.037\mathrm{kpc^{-2}Myr^{-1}}$, about three times higher than for the MWSC sample.
As we note above, the observed CAMF generally sits at lower ages compared to the MWSC data. Likewise, the CAMF of the best-fitting model has high-age edge, density maximum and violent relaxation mass loss at higher ages compared to observation.
We used steps of 0.05 dex to find that a shift of the model to lower ages by 0.8 dex minimises the KLD between observed CAMF and error-smoothed model CAMF to $D_{\mathrm{KL}}=0.113$ from $D_{\mathrm{KL}}=0.666$ without an age shift. Such a shift in cluster age is equivalent to rescaling the parameters $\beta$, $c$, and $t_{\mathrm{V}}$, which respectively govern cluster formation rate, lifetimes, and the violent relaxation timescale.
@@
We finally used the KLD and the above error model to perform a model fit using MCMC sampling as we did for the MWSC sample. For simplicity, we reused the same parameter constraints as they fit also this sample. For the Gaussian priors, we adopted the same means as a baseline, but used twice the standard deviations to account for the Gaussian priors being specific to the MWSC sample and avoid unduly constraining the parameter space.
The MCMC took 221000 steps with an average acceptance rate of 0.161 until convergence. With a burn-in of 4404 samples and a thinning factor of 528, the final posterior sample contains 21320 parameter sets. The best model has a KLD of $D_{\mathrm{KL}}=0.034$, which represents a significant improvement compared to the age-shifted best-fitting MWSC model. The rightmost column of Table 4 contains a list of this model’s parameters with uncertainties derived from the posterior distribution.
@@
Indeed the best model, with its CAMF shown in Fig. 12, does not feature a break in the cluster lifetime-mass relation and has a high value of $a_{1}$ at 0.538, resulting in slow dissolution at the end of clusters’ lifetimes and a high number of clusters with masses below $40\mathrm{M}_{\odot}$. The Schechter cutoff mass at $\log m_{\mathrm{S}}/\mathrm{M}_{\odot}=4.743$ corresponds to lifetimes close to the age limit and seems mostly sufficient to reproduce it.
@@
The model contours in the left panel of Fig. 12 appear to match the observed CAMF fairly well. Synthetic cluster samples drawn from the model contain on average $2447\pm 41$ clusters for a total surface density of $\Sigma_{\mu>1.6}=257.22\pm 0.50\mathrm{kpc}^{-2}$ above the mass limit, which appears consistent with observations within the stochastic variation. However, all 5000 synthetic samples we drew have lower values of the KLD than the observed sample, indicating the presence of systematic differences between observation and model that cannot be explained by stochastic noise.
@@ 7 Summary and conclusions
In this paper we investigated the parameter space of a theoretical model for the cluster age-mass function (CAMF), which was introduced in Just et al. (2023), where a base model was presented. The model was constructed with a constant cluster formation rate (CFR), a two-power law cluster initial mass function (CIMF) with an exponential Schechter cutoff, and a cluster bound mass function $m(M,t)$ describing the cluster mass loss as function of initial mass and age. The cluster mass evolution is adapted to the models of Shukirgaliyev et al. (2017, 2018) based on a centrally concentrated star formation efficiency.
@@
A posterior analysis by random sampling of the models was used to determine whether a configuration similar to the observed cluster sample is a likely configuration of such a model sample. For this, the different statistics as well as the total cluster number were used to quantify deviations between models and their samples, and assess whether the observed cluster sample can be differentiated from a model sample based on the observed values of the different statistics. The base model, published in Just et al. (2023), can be differentiated from observations by $\chi^{2}$, KLD, and total cluster number, but not by the maximum likelihood method. The models returned from $\chi^{2}$ and maximum likelihood can be differentiated less easily, but also show deviations in at least two of the statistics considered. The model returned from the KLD in particular is best able to reproduce the observed cluster number and can only be differentiated from observations by the maximum likelihood method.
Using principal component analysis, we derived the ellipsoidal shape of the 11-dimensional posterior distribution in parameter space for the different MCMC runs. The semi-major axes of the ellipsoids (given by the square root of the normalised covariance matrix eigenvalues) vary by a factor of four, for example from ${\sqrt{\lambda_{0}}=0.341}$ to ${\sqrt{\lambda_{10}}=1.486}$ for the $D_{\mathrm{KL}}$-fit (see also Fig. 5). These eigenvalues give the variance of the parameter vector in the direction of the eigenvector relative to the individual parameter variances. An eigenvalue smaller than 1 corresponds to a combination of parameters that is more tightly constrained than the parameters are individually, while an eigenvalue larger than 1 corresponds to a combination that is constrained less so. The eigenvectors corresponding to both the smallest and largest eigenvalues are similar for the posteriors of the different runs, indicating that the different fit statistics realise the same constraints imposed by the observed CAMF. The model parameters are very well constrained in two dimensions fixing the upper age limit and the position of the maximum of the CAMF and badly constrained in one dimension corresponding to low-mass end of the distribution.
@@
The base model was constructed to reproduce the mass functions in four age bins (Just et al. 2023). The focus was on the younger ages $\log t/\mathrm{yr}<8.3$ and the maximum of the number surface density. The best model resulting from the $D_{\mathrm{KL}}$-fit, which we adopted as our best-fitting model, provides a similarly good representation of these features of the CAMF and a significantly improved model for the old and the high-mass regions.
@@
However, there are two main challenges in inferring model parameters and selecting best-fit models we identified over the course of this work. Both challenges are expected to not be specific to the model functions and the cluster data we used, but intrinsic to the problem of modelling the evolution of a cluster system. We identified these first in the context of the MWSC dataset and found them to also be relevant when using the Hunt & Reffert (2023, 2024) catalogue. The first challenge lies in the strongly correlated nature of model parameter space, which makes constraining individual parameters difficult. For instance, the rates of cluster formation and dissolution must be balanced such that the model reproduces the observed cluster content, but individually each of these rates are badly constrained. A possible remedy for this problem may lie in better and more sophisticated constraints on cluster mass evolution from prior physics considerations such as N-body simulations. The second challenge lies in the measuring of observation-model correspondence, where uncertainties in the measurements of observed cluster properties, the completeness limits of the observed cluster sample and the varying cluster density across the age-mass plane must be taken into consideration. In particular, the resolution of the observed CAMF in the age-mass plane is limited by both sample size and measurement uncertainties, which can make approaches that work without explicit construction of the observed CAMF, such as our expected log-likelihood approach, more attractive. We demonstrate the use of forward modelling to check the behaviour of a fit statistic for a given model, and discuss how a model fitting procedure can be limited by the fit statistic used. More work is needed to identify efficient statistics for measuring the correspondence of a CAMF model to a cluster sample, especially in view of the extensive, high-quality cluster catalogues based on Gaia DR3 (Gaia Collaboration et al. 2023) that are currently emerging (e.g. Hunt & Reffert 2023).
@@ Appendix A Rapid mass evolution in the observed CAMF using linear age binning
We argue in Sect. 2 that the constant high-mass edge for $\tau<7.8$ in the observed CAMF is evidence for rapid mass evolution during the violent relaxation phase. One way to more accessibly visualise this is to remove the effect of logarithmic age binning by normalising using linear bin sizes in age instead (as also used in Fig. 8 to visualise the factorisation of the CAMF in initial-mass).
@@ Appendix B The modelled CAF
As is discussed in Sect. 1, the CAF is typically of interest in the context of fitting cluster evolution models. While our 2D approach does not consider or fit the CAF explicitly, correspondence between observed and modelled CAF should nonetheless arise by construction.
For our model, the CAF $\eta(t)$, giving the cluster surface density per age interval, can be computed at some age $t$ by integrating the CAMF over all masses. Recalling Eqs. (10) and (46), with $M_{0}(t)$ as inverse of $t_{\mathrm{Cl}}(M)$, we find